OpenStudy (anonymous):

I am starting on my quiz due tomorrow and need to know if this is right. sqrt(36/49)=.857

OpenStudy (anonymous):

6/7 = 0.866

OpenStudy (anonymous):

So do I round them ?

OpenStudy (anonymous):

It depends on what kind of calculator you're using! Let me check mine.

OpenStudy (anonymous):

Mine said it was0.857.

OpenStudy (anonymous):

Yep, mine gets the same as yours!

OpenStudy (anonymous):

Do you have any other question?

OpenStudy (anonymous):

Thanks for sticking around to answer my question instead of just giving me the answer and leaivng.

OpenStudy (anonymous):

I have several problems that I am working on...I don't get this algebra stuff at all.

OpenStudy (anonymous):

Even I'm leaving, all you do is typing on the message, we are notified and fly back in no time :)

OpenStudy (anonymous):

Go ahead, ask as much as you can :)

OpenStudy (anonymous):

:) Thanks a bunch! You have been super helpful.

OpenStudy (anonymous):

17sqrt(48)+8sqrt(27)

OpenStudy (anonymous):

I learn this manner from most of OS masters :)

OpenStudy (anonymous):

Well thank you!

OpenStudy (anonymous):

17√(48) + 8√(27) = 17 √ (16 * 3 ) + 8 √ 9 * 3 = 17 * 4 √3 + 8 * 3 √3 = 68√3 + 24√3 = 92√3 ( 159.3)

OpenStudy (anonymous):

BTW, are you sure that you're allowed to use calculator in exam?

OpenStudy (anonymous):

and thank you for showing me the work also...it makes it much more understandable! And yes I can use a calculator, its just a quiz in mymathlab, I am an online student.

OpenStudy (anonymous):

If not, then the result is 92√3 !

OpenStudy (anonymous):

It didn't specify whether or not to use the sqrt answer or not. so I don't know.

OpenStudy (anonymous):

Since the purpose of Math at your level is learning how to apply exponent, I don't think ultilize calculator will be helpful!

OpenStudy (anonymous):

This question says to find the following and assume that a and b can represent any real number. \[\sqrt[432]{a+b}^{432}\]

OpenStudy (anonymous):

So your suggestion is to not use a calculator and give the real answer?

OpenStudy (anonymous):

I just want to make sure you benefit from your study!

OpenStudy (anonymous):

I understand that....I like seeing the work b/c it helps me understand it some, but I can't get it down in my head how this stuff works. My quiz scores have been really low b/c I just can't get it.

OpenStudy (anonymous):

(a + b)^(432/432) = (a + b)^1 = a + b

OpenStudy (anonymous):

When I solving the problem, whatever I'm thinking I'm writing it down.

OpenStudy (anonymous):

I have resorted to using flash cards with example problems on them. That doesn't even help me most of the time.

OpenStudy (anonymous):

With math and science, the only way to learn effectively is practice like you doing now!

OpenStudy (anonymous):

Anyways, thank you so much for your help. I am gonna try to work some of this on my own. I still have 20 questions to go.

OpenStudy (anonymous):

Work on them! Just post then I explain it right away :)

OpenStudy (anonymous):

Thank you...I will keep OS pulled up.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Ok for this problem \[\sqrt[4]{16}*\sqrt[4]{32}=4*\sqrt[4]{2}\]

OpenStudy (anonymous):

Mulitiply and simplify by factoring.

OpenStudy (anonymous):

that is the answer I have...

OpenStudy (anonymous):

= (2^4) ^1/4 = 2 * 2 √2 = 4√2

OpenStudy (anonymous):

I mean fourth root of 2

OpenStudy (anonymous):

Ok so the 4* doesn't belong in that answer?

OpenStudy (anonymous):

The answer is

OpenStudy (anonymous):

your answer's right!

OpenStudy (anonymous):

^4sqrt(2)

OpenStudy (anonymous):

Ok awesome...thank you. I am sure I will be back again!

OpenStudy (anonymous):

I'm just lazy to write decently like you do :")

OpenStudy (anonymous):

Its ok...

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Here is one...